joshkconnell:

The sum of two consecutive terms in the arithmetic sequence 4, 9, 14, 19,... is 503; find these two terms. the first consecutive term of the arithmetic sequence is __? the second __?

darkknight:

first you would like to find the eqn for the arithmetic sequence. As a hint plus 5 every term. Afterwards you want to find the 2 terms that add to 503

darkknight:

so lets write the eqn out plus 4 every term so we have a formula \[F = a +(n-1)d\] the sequence starts out with 4, so when 1 = 1 then F = 4 So we know that a = 4

darkknight:

continuing from here, we need to find d when n = 2, F =9 So 9 = 4 + (2-1)(d) 9 = 4 + d d = 5

darkknight:

now lets test if this eqn works \[F= 4+(n-1)5\]

darkknight:

plug in 3 for n and see if you get 14